OpenStudy (anonymous):

What is the location of this parabola's focal point? y = (1/4)(x – 400)2

OpenStudy (dumbcow):

p is distance from vertex to focus \[p = \frac{1}{4a}\] where \[y = a(x-h)^{2} +k\] so plug in "a" to find "p" then add that to coord of vertex

OpenStudy (anonymous):

but what is a @dumbcow

OpenStudy (dumbcow):

its the coefficient in front of (x-400)^2 term hence the general parabola equation \[y = a(x-h)^{2} +k\]

OpenStudy (anonymous):

y=1/4(x-h)^2+k then?

OpenStudy (dumbcow):

no sorry im confusing you ... im giving the general form...your equation is \[y = \frac{1}{4}(x-400)^{2}\] these numbers correspond to coefficient in general form a = 1/4 h = 400 k=0

OpenStudy (anonymous):

Oh okay @dumbcow

OpenStudy (anonymous):

I need to have the answer in an ordered pair, though @dumbcow

OpenStudy (dumbcow):

focus is always directly above or below the vertex ... if its a function of "x" so "x-coord" is same as vertex

OpenStudy (dumbcow):

vertex: (h,k) thus focus: (h, k+p)

OpenStudy (anonymous):

@dumbcow thank you

OpenStudy (dumbcow):

your welcome